# GDPR Elements, and Future Directions

### v3

@@ V-B4 Using blockchain as a multimodal biometric authentication system
Blockchain helps the multimodal biometric system by providing a decentralized, secure, and privacy-preserving mechanism to manage access to the distributed stored encrypted biometric templates. Addresses the vulnerabilities of centralized systems by eliminating the single point of failure, reducing the risk of data breaches, improving system availability, and empowering users with control over their biometric data through smart contracts. The results of each modality can be published and linked on the blockchain. As a result, several applications can authenticate the same individual simultaneously. In addition, most of the time, the underlying biometric hardware that captures user biometrics is not made or maintained by the same party. Therefore, publishing the authentication result for each modality on the blockchain helps to eliminate inconsistencies between the underlying infrastructure of each hardware. These results can be linked using an identifier, for example, using DIDs. Therefore, it is possible to link the biometric modalities of individuals using DIDs on the blockchain. Although standardization and risk mitigation efforts, such as DIDs, are appreciated, it should be kept in mind that DID documents can still be linked to individuals, and hence they are considered personal data under legal rules like the GDPR. Another practice is the distributed storage with IPFS. To this end, a recent study [130] stored decentralized fuzzy vaults of facial biometric images and dorsal hand images on IPFS and their address on the Ethereum Virtual Machine (EVM)-based blockchain network. Once the addresses of the fuzzy vaults are recorded on the blockchain, they become immutable, making it extremely difficult for malicious attackers to manipulate or delete this crucial information without consensus from the network participants.
@@ V-C Suggestions to Mitigate Legal Risks
First of all, we suggest that biometric data and any other personal data, such as name, nationality, etc., should not be stored on-chain until further research confirms that protective measures are sufficient. [131, 105].
Furthermore, to mitigate data protection risks, the objective of processing operations must be clearly stated and compared to the applicable provisions outlined above to comply with GDPR. For instance, in cases where the purpose of the processing is in the public interest, e.g., electronic voting system, public health, or processing personal data on-chain to facilitate compliance with the COVID-19 restrictions at the borders, the exemptions to the right to erasure (Article 17 GDPR) could be relied on.
@@
V-D Recommendations for Interdisciplinary Research
To improve the understanding and implementation of data protection by design, it is crucial, first of all, to establish a standard and harmonized understanding of the key terminology between legal researchers and computer scientists. This recommendation explicitly regards the term anonymity as explained, which is often misunderstood by the scientific community. In addition, privacy-preserving technologies should not only be understood as a means of improving data security. Although highly important, data security is only one aspect of the data protection framework. Thus, when developing privacy-preserving technologies, technology should incorporate other data protection principles enshrined under the GDPR, such as purpose limitation, lawfulness, data minimization, storage limitation, and accountability. Additionally, data protection analysis should not only focus on generic blockchain architectures, but also on additional technologies, as each project utilizes different complementary technologies to increase the efficiency and security of their systems. We call for specific guidance to mitigate these risks, taking into account the sensitive nature of biometric data.
@@ VI Conclusions
In this survey, we investigated the advantages and shortcomings of the mutual integration of blockchain and biometric recognition systems. The goal of this survey is to clear the path for this research line to further develop and to clarify some of the legal requirements while considering the technical status and advancements of this technology. To sum up, this combination can be favorable for biometrics as it provides a platform for PKI, distributed identity management, federated learning, and multimodal biometrics. On the other hand, we found that it contains risks for real-time applications or when used as a storage space for sensitive information such as biometric data. Nevertheless, with further advances in blockchain technology, some of these conditions may change. From a legal point of view, one shall take into account the sensitive nature of biometric data and the specific emerging risks out of this combination, such as accountability. It is important when experimenting with biometric data and blockchain, researchers keep both raw biometric data (images) and templates off-chain. Moreover, the purpose of the processing should be clarified for each granular processing operation, and the risks should be assessed in order to apply additional safeguards for a future-proof data protection design. Better precision of the purpose of the processing can also help understand whether the exemptions to the right to be forgotten (Article 17 GDPR) apply to that processing. Finally, considering the current literature and the insights of the authors, we provided some recommendations and future directions for combining blockchain and biometrics, such as achieving a harmonized language between legal and technical research.

### v4

@@ V-B4 Using blockchain as a multimodal biometric authentication system
Blockchain helps the multimodal biometric system by providing a decentralized, secure, and privacy-preserving mechanism to manage access to the distributed stored encrypted biometric templates. It addresses the vulnerabilities of centralized systems by eliminating single points of failure, reducing the risk of data breaches, improving system availability, and empowering users with control over their biometric data through smart contracts. The results of each modality can be published and linked on the blockchain. As a result, several applications can authenticate the same individual simultaneously. In addition, most of the time, the underlying biometric hardware that captures user biometrics is not made or maintained by the same party. Therefore, publishing the authentication result for each modality on the blockchain helps to eliminate inconsistencies between the underlying infrastructure of each hardware. These results can be linked using an identifier, for example, using DIDs. Therefore, it is possible to link the biometric modalities of individuals using DIDs on the blockchain. Although standardization and risk mitigation efforts, such as DIDs, are appreciated, it should be kept in mind that DID documents can still be linked to individuals, and hence they are considered personal data under legal rules like the GDPR. Another practice is the distributed storage with IPFS. To this end, a recent study [138] stored decentralized fuzzy vaults of facial biometric images and dorsal hand images on IPFS and their address on the Ethereum Virtual Machine (EVM)-based blockchain network. Once the addresses of the fuzzy vaults are recorded on the blockchain, they become immutable, making it extremely difficult for malicious attackers to manipulate or delete this crucial information without consensus from the network participants.
@@ V-C Suggestions to Mitigate Legal Risks
First of all, we suggest that complete biometric data and any other personal data, such as name, nationality, etc., should not be stored on-chain until further research confirms that protective measures are sufficient [139, 112].
Furthermore, to mitigate data protection risks, the objective of processing operations must be clearly stated and compared to the applicable provisions outlined above to comply with GDPR. For instance, in cases where the purpose of the processing is in the public interest (e.g., electronic voting system, public health, or processing personal data on-chain to facilitate compliance with the COVID-19 restrictions at the borders), the exemptions to the right to erasure (Article 17 GDPR) could be relied on.
@@ VI Conclusions
This survey examined the integration of blockchain and biometric recognition systems, highlighting both their potentials and limitations. From a technical standpoint, blockchain offers a robust foundation for enhancing the integrity, auditability, and interoperability of biometric systems. Its decentralized nature enables trust without reliance on a central authority, and recent frameworks demonstrate its capacity to support secure identity management and transparent verification. While real-time biometric applications and on-chain storage of sensitive data remain challenging, notable progress has been achieved through off-chain storage with on-chain proofs, supported by distributed file systems such as IPFS and permissioned frameworks like Hyperledger. Moreover, the emergence of high-throughput blockchain networks holds great promise for overcoming latency and scalability constraints, paving the way for more responsive and efficient biometric systems. From a legal point of view, one shall take into account the sensitive nature of biometric data and the specific emerging risks out of this combination, such as accountability. It is important when experimenting with biometric data and blockchain, researchers keep both raw biometric data (images) and templates off- chain. Moreover, the purpose of the processing should be clarified for each granular processing operation, and the risks should be assessed in order to apply additional safeguards for a future-proof data protection design. Better precision of the purpose of the processing can also help understand whether the exemptions to the right to be forgotten (Article 17 GDPR) apply to that processing. Finally, based on the reviewed literature and our analysis, we highlight several recommendations and research directions for future studies, including the development of GDPR-compliant on-chain designs, the exploration of high-throughput blockchain platforms for near real-time verification, and the investigation of blockchain as a decentralized infrastructure for biometric-based PKI.